OpenStudy (anonymous):

Simplify by removing the factors of 1. 10y^2-40/25y^2-100 please show the steps

OpenStudy (anonymous):

simplifying the numerator we get 10y^2-40=10(y^2-4)=10(y+2)(y-2)

OpenStudy (anonymous):

simplifying the denominator we get 25y^2-100=25(y^2-4)=25(y+2)(y-2)

OpenStudy (anonymous):

our new expression should look like this 10(y+2)(y-2)/25(y+2)(y-2)

OpenStudy (anonymous):

so it is going to be (y+5)(y-5)

OpenStudy (anonymous):

now we simply cancel (y+2)(y-2) since it is a coefficient on bothe the numerator and the denominator to get 10/25=2/5 as our answer.

OpenStudy (anonymous):

so it is going to be in a fraction form

OpenStudy (anonymous):

yes your answer should be a fraction